Solve using the Master Method

T(n) = 3T(n/2) + n
Answer #1

Solution:

Masters Theorem:

if the function is of the form

T(n) = at(n/6) + f(n),a >=1,6 > 1

then

if f(n) = O(n) and c < logga then T(n) = O(nlogia) ---------------(eq 1)

if f(n) = O(n) and c=logya then Tn) = O(nlogn)) ------------(eq 2)

if f(n) = O(n) and c> logga then T(n) = O(fn)) --------------------(eq 3)

So,let the given equation be,

T(n) = 3T (n/2) +n

in the above equation we can identify the values of

a=3

b=2

f(n) = n

c=1`

Now,

log23 = 1.58 and c=1

therefore we can conclude that,

c<log23

therefore from eq 1 of masters theorem we have,

T(n)=O(n^{log_{b}a})

T(n)=O(n^{log_{2}3})

T(n)=O(n^{1.58})